OpenStudy (anonymous):

Match Term Definition positive thirty−two minus positive seven A) negative thirty−nine positive thirty−two minus negative seven B) positive thirty−nine negative thirty−two minus positive seven C) negative twenty−five negative thirty−two minus negative seven D) positive twenty−five

OpenStudy (anonymous):

@Chineseboy15

OpenStudy (anonymous):

30+(2-7) As 2-7=-5 So 30+(-5) =30-5 =25 The answer is D

OpenStudy (anonymous):

Try to solve other three questions yourself.